OpenStudy (anonymous):

A container has 50 electronic components, of which 10 are defective. If 6 components are drawn at random from the container, what is the probability that at least 4 are good? 0.26 0.42 0.75 0.91 1.00

OpenStudy (anonymous):

A If 4 are good, 2 must be defective. ----- Binomial with n = 6 and p(defective) = 1/5 ---- P(x = 2) = 6C2*(1/5)^2*(4/5)^4 = binompdf(6,1/5,2) = 0.2458 =

OpenStudy (anonymous):

thanks for helping me out i suck in geometry